# An Overview of Prototype Formulations for Interpretable Deep Learning

### v3

@@ Abstract
Prototypical part networks offer interpretable alternatives to black-box deep learning models. However, many of these networks rely on Euclidean prototypes, which may limit their flexibility. This work provides a comprehensive overview of various prototype formulations. Experiments conducted on the CUB-200-2011, Stanford Cars, and Oxford Flowers datasets demonstrate the effectiveness and versatility of these different formulations.
@@ 1 Introduction
Deep Learning has achieved high accuracy in many computer vision tasks. However, the decision-making processes of these models lack transparency and interpretability, making deployment in safety-critical areas challenging. Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of such deep models and increase trust in their decisions.
XAI methods can be broadly divided into two categories: First, Post-Hoc Methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) or GradCAM (Selvaraju et al., 2017) offer explanations for predictions without requiring retraining. While applicable in many scenarios, post-hoc methods may not actually align with the models’ decision-making processes, potentially leading to interpretations that are not entirely faithful (Rudin, 2019). Second, inherently interpretable methods provide built-in, case-based reasoning processes. For instance, small decision trees are inherently interpretable because their reasoning can be easily understood as a series of if-else statements (Molnar, 2020). However, they are constrained in their representational power.
@@
Deep Prototype Learning Architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (e.g., Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning models through a prototype layer. Each neuron in this layer represents a prototype, storing a latent feature vector. The model’s predictions are based on the distances between sample features and prototype parameters, for example, by computing the $L_{2}$-distance. This makes prototype learning essentially a clustering task in latent space (Zhou et al., 2022). Different prototype formulations, such as prototypes using the cosine similarity, induce a hyperspherical structure to the latent space with performance advantages to classification (Mettes et al., 2019). Probabilistic formulations such as Gaussian prototypes enable further downstream tasks relying on probability values such as Bayesian approaches.
With the increased popularity of prototypical part models, there has been a corresponding rise in the number of architectural choices. However, to the best of our knowledge, there is no comprehensive overview of the design choices and prototype formulations for learning interpretable prototypes and their effect on predictive performance.
Table 1: Overview of 14 related models and their prototype learning configurations. 8/14 models use Euclidean prototypes.
@@
Prototype Decoder Euclidean Entire Image Class Exclusive FCL Li et al. (2018)
ProtoPNet Euclidean Patch Class Exclusive FCL Chen et al. (2019)
Def. ProtoPNet Cosine Spatial Arrangement Class Exclusive FCL Donnelly et al. (2021)
ProtoPShare Euclidean Patch Merged after Training FCL Rymarczyk et al. (2020)
ProtoPool Euclidean Patch Pooling Shared FCL Rymarczyk et al. (2022)
TesNet Grassman Patch Class Exclusive FCL Wang et al. (2021)
ProtoTree Euclidean Patch Shared Decision Tree Nauta et al. (2021)
ProtoKNN Cosine Patch Class Exclusive KNN Clf Ukai et al. (2023)
PIPNet Cosine Patch Shared FCL Nauta et al. (2023)
LucidPPN Euclidean Separate Color & Texture Class Exclusive Branch Aggregation Pach et al. (2024)
ProtoSeg Cosine Patch Class Exclusive FCL Zhou et al. (2022)
ProtoGMM Gaussian Patch Class Exclusive FCL Moradinasab et al. (2024)
MGProto Gaussian Patch Class Exclusive Bayesian Likelihood Wang et al. (2024b)
ProtoPFormer Euclidean Transformer Token Class Exclusive Branch Aggregration Xue et al. (2024)
This paper explores different prototype formulations. Its contributions are as follows:
Prototype Formulation Overview: A comprehensive overview of different, learnable prototype definitions, including both point based and probabilistic prototypes as shown in Figure 1.
HyperPG: A new probabilistic prototype representation with learned parameters anchor $\alpha$, mean $\mu$ and variance $\sigma^{2}$. This representation models a Gaussian distribution over cosine similarities, thereby projecting a Gaussian distribution on the surface of a hypersphere. HyperPG’s similarity score is based on the Gaussian’s probability density function and adapts its size through a learned standard deviation.
Benchmarking Protoypes: Extensive image classification experiments based on prototypical part network architectures like ProtoPNet on the CUB-200-2011 (Wah et al., 2011), Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets.
@@ 2 Related Work
All these approaches share the use of a linear classification layer. However, ProtoTree (Nauta et al., 2021) builds a Decision Tree on the learned Euclidean prototypes and ProtoKNN (Ukai et al., 2023) proposes to use a $k$-nearest-neighbor classifier with cosine prototype similarities.
@@
Zhou et al. (2022) propose to view prototype learning as a clustering task in the network’s latent space for image segmentation. In this setting, the clustering can be done outside of the network’s backpropagation path with an external clustering algorithm, so-called non-learnable prototypes. Later work such as ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2024b) build on this idea and model the prototypes as Gaussian distributions.
ProtoPFormer (Xue et al., 2024) adapts Euclidean prototypical part networks to Transformer based architectures by using a VisionTransformer (ViT, Dosovitskiy, 2020) backbone. To avoid the distraction problem due to the global attention mechanism of ViT architectures, ProtoPFormer uses a hybrid approach with global and local prototypes. The local prototypes are encouraged to focus on different parts of the image by modeling the spatial prototype activations as a 2D Gaussian.
Table 1 presents a systematic overview of the discussed works. The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019), but only few prototype learning works have used them so far (Donnelly et al., 2021; Zhou et al., 2022; Ukai et al., 2023). There is also a lack for probabilistic prototypes in a hyperspherical space. We introduce HyperPG, which combines the predictive performance gains from the cosine similarity with the probabilistic nature of Gaussian distributions.
@@ 3 Overview: Prototype Formulations
Figure 2: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$. Euclidean prototypes compute the $L_{2}$ distance between two points in latent space. Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere. Gaussian prototypes model a Gaussian distribution in Euclidean space and compute the probability density function (PDF). HyperPG prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
@@ 3.1 Point Based Prototypes
The prediction is computed by comparing each prototype ${\bm{p}}$ to the latent feature map ${\bm{z}}$. For simplicity’s sake lets assume the spatial dimensions $\rho=\zeta=1$. The following equations can be adapted for higher spatial dimensions by summing over the height and width $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each patch of the latent map.
@@ 3.1.2 Cosine Prototypes
which is based on the angle between two normalized vectors of unit length. By normalizing $D$ dimensional vectors to unit length, they are projected onto the surface of a $D$ dimensional hypershere. The cosine similarity is defined on the interval $[-1,1]$ and measures: $1$ for two vectors pointing in the same direction, 00 for orthogonal vectors, and $-1$ for vectors pointing in opposite directions. Without the normalization to unit length, we can define a scaled dot-product prototype inspired by the Attention mechanism (Vaswani et al., 2017). However, this similarity measure is not bounded to any interval, as the vector length is proportionally preserved. For more details see Appendix C. Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing two vectors directly.
@@ 3.2.1 Gaussian Prototypes
It is worth noting that the term inside the exponent is based on the $L_{2}$ distance between the latent vector ${\bm{z}}$ and the mean $\bm{{\mu}}$. This formulation as a PDF has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space. B) Prototypes can adapt their shape using a full covariance matrix, allowing different variances along various feature dimensions, offering more flexibility in shaping the latent space. However, learning a full covariance matrix increases computational requirements, especially with EM clustering.
@@
Figure 3: HyperPG prototypes learn a distribution of cosine similarities. They are parameterized by a learnable anchor vector $\bm{{\alpha}}$, scalar mean $\mu$ and scalar variance $\sigma^{2}$. HyperPG projects a Gaussian distribution of cosine similarities on the surface of a hypersphere, resulting in ring shaped activation patterns around the anchor vector.
@@ 3.2.2 Gaussian Prototypes on the Hypersphere - HyperPG
Just as Gaussian prototypes provide a probabilistic formulation for Euclidean prototypes, we aim to develop a probabilistic formulation for hyperspherical prototypes based on the cosine similarity. We name this new formulation Prototypical Gaussians on the Hypersphere (HyperPG). HyperPG prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$. HyperPG prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$. Because the cosine similarity is bounded to $[-1,1]$, HyperPG’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds. Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function. Then, HyperPG’s similarity measure based on the truncated Gaussian distribution is defined as
@@
Figure 3 illustrates the activations of HyperPG’s similarity function on the surface of a $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, fixed std $\sigma=0.1$ and various mean values $\mu\in[0,1]$. The anchor $\bm{{\alpha}}$ defines a prototypical direction vector in latent space $\mathcal{Z}$, similar to other hyperspherical prototypes, and is visualized as a red arrow. The learned Gaussian distribution of cosine similarities is projected onto the hypersphere’s surface with std $\sigma$ governing the spread of the distribution, and mean $\mu$ the expected distance to the anchor $\bm{{\alpha}}$. For $\mu=1$, the distribution centers around the anchor, as the cosine similarity is $1$ if two vectors point in the same direction. This corresponds to the von Mises-Fisher (vMF) distribution (Appendix B). For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is $-1$ for vectors pointing in opposite directions.
For values of $11>\mu>-$ 1, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$. This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction. Imagining the interpolation between $\mu=1$ and $\mu=-1$, the probability mass moves from one pole of the hypersphere to the other, stretching like a rubber band over the surface. For $\mu=0$, the expected cosine similarity indicates that vectors with the highest activation are orthogonal to the anchor $\bm{{\alpha}}$. Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation.
@@
HyperPG can be easily adapted to other probability distributions with additional desirable properties. Possible candidate distributions are elaborated on in Appendix B. Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPG idea transferable to other manifolds and applications outside of prototype learning.
4 Training

### v4

@@ Abstract
Prototypical part networks offer interpretable alternatives to black-box deep learning models by learning visual prototypes for classification. This work provides a comprehensive analysis of prototype formulations, comparing point-based and probabilistic approaches in both Euclidean and hyperspherical latent spaces. We introduce HyperPG, a probabilistic prototype representation using Gaussian distributions on hyperspheres. Experiments on CUB-200-2011, Stanford Cars, and Oxford Flowers datasets show that hyperspherical prototypes outperform standard Euclidean formulations. Critically, hyperspherical prototypes maintain competitive performance under simplified training schemes, while Euclidean prototypes require extensive hyperparameter tuning.
@@ 1 Introduction
Deep Learning has achieved high accuracy in many computer vision tasks. However, the decision-making processes of these models lack transparency and interpretability, making deployment in safety-critical areas challenging. Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of models and increase trust in their decisions.
XAI methods can be broadly divided into two categories: First, post-hoc methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017) or GradCAM (Selvaraju et al., 2017), which explain model predictions without retraining. While broadly applicable, post-hoc methods may not align with the models’ actual decision-making processes (Rudin, 2019). Second, inherently interpretable methods provide built-in, case-based reasoning. For instance, small decision trees are inherently interpretable through their if-else structure (Molnar, 2020), but are limited in representational power.
@@
Deep prototype learning architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning through a prototype layer. Each prototype neuron stores a latent feature vector, with predictions based on distances between sample features and prototypes. This makes prototype learning a clustering task in latent space (Zhou et al., 2022). Different formulations induce different latent structures: cosine similarity induces hyperspherical structure with classification advantages (Mettes et al., 2019), while probabilistic formulations like Gaussian prototypes enable Bayesian downstream tasks.
Table 1: Overview of 15 related models and their prototype learning configurations.
@@
Prototype Decoder Euclidean Entire Image Class Exclusive FCL (Li et al., 2018)
ProtoPNet Euclidean Patch Class Exclusive FCL (Chen et al., 2019)
Def. ProtoPNet Cosine Spatial Arrangement Class Exclusive FCL (Donnelly et al., 2021)
ProtoPShare Euclidean Patch Merged after Training FCL (Rymarczyk et al., 2020)
ProtoPool Euclidean Patch Pooling Shared FCL (Rymarczyk et al., 2022)
TesNet Grassman Patch Class Exclusive FCL (Wang et al., 2021)
ProtoTree Euclidean Patch Shared Decision Tree (Nauta et al., 2021)
ProtoKNN Cosine Patch Class Exclusive KNN Clf (Ukai et al., 2023)
PIPNet Cosine Patch Shared FCL (Nauta et al., 2023)
ST-ProtoPNet Cosine Patch Class Exclusive Branch Aggregation (Wang et al., 2023)
LucidPPN Euclidean Separate Color & Texture Class Exclusive Branch Aggregation (Pach et al., 2024)
ProtoSeg Cosine Patch Class Exclusive FCL (Zhou et al., 2022)
ProtoGMM Gaussian Patch Class Exclusive FCL (Moradinasab et al., 2024)
MGProto Gaussian Patch Class Exclusive Bayesian Likelihood (Wang et al., 2024b)
ProtoPFormer Euclidean Transformer Token Class Exclusive Branch Aggregration (Xue et al., 2024)
The increased popularity of prototypical part models increases architectural choices, yet there is no comprehensive overview of prototype formulations and their effect on predictive performance. This paper explores different prototype formulations with the following contributions:
  • Prototype Formulation Overview: A comprehensive overview of different learnable prototype definitions, including both point-based and probabilistic prototypes as shown in Figure 1.
  • HyperPG: A probabilistic prototype representation modeling Gaussian distributions over cosine similarities with learned anchor $\alpha$, mean $\mu$, and variance $\sigma^{2}$, projecting Gaussians onto hypersphere surfaces.
  • Benchmarking and Robustness Analysis: Image classification experiments on CUB-200-2011 (Wah et al., 2011), Stanford Cars (Krause et al., 2013), and Oxford Flowers (Nilsback and Zisserman, 2008) datasets, evaluating both full optimization and simplified training procedures to reveal differences in hyperparameter sensitivity. Our code is publicly available on Github [^1] https://github.com/LiXiling/prob-proto.
@@ 2 Related Work
Figure 2: Illustration of similarity computation between prototype $\bm{p}$} and latent vector $\bm{z}$} for different formulations. Euclidean: $L_{2}$ distance in latent space. Hyperspherical: cosine similarity of normalized vectors (angle on hypersphere). Gaussian: PDF of Gaussian distribution in Euclidean space. HyperPG: PDF of Gaussian distribution over cosine similarities (Gaussian on hypersphere surface).
All these approaches share the use of a linear classification layer. However, ProtoTree (Nauta et al., 2021) builds a Decision Tree on the learned Euclidean prototypes and ProtoKNN (Ukai et al., 2023) proposes to use a $k$-nearest-neighbor classifier with cosine prototype similarities. ST-ProtoPNet (Wang et al., 2023) learn the prototypes close to the decision boundary, similar to support vectors.
@@
“Rethinking Semantic Segmentation” (Zhou et al., 2022) view prototype learning as a clustering task in the network’s latent space. The clustering is done outside of the network’s backpropagation path with an external clustering algorithm, so-called non-learnable prototypes. Later work such as ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2024b) build on this idea and model the prototypes as Gaussian distributions.
ProtoPFormer (Xue et al., 2024) adapts Euclidean ProtoPNets to Transformers through a ViT (Dosovitskiy, 2020) backbone. To avoid the distraction problem due to the global attention mechanism of ViTs, ProtoPFormer uses a hybrid approach with global and local prototypes. The local prototypes are encouraged to focus on different parts of the image by modeling the spatial prototype activations as a 2D Gaussian.
Table 1 presents a systematic overview, where the cosine similarity performs well in classification tasks (Mettes et al., 2019), but only few prototype learning works have used them so far (Donnelly et al., 2021; Zhou et al., 2022; Ukai et al., 2023). There is a lack for probabilistic prototypes in a hyperspherical space. We introduce HyperPG, that combines the performance gains from the cosine similarity with the probabilistic nature of Gaussian distributions.
@@ 3.1 Point Based Prototypes
The prediction is done by comparing each prototype $\bm{p}$} to the latent feature map $\bm{z}$}. For simplicity, lets assume the spatial dimensions $\rho=\zeta=1$. The following equations can be adapted for higher spatial dimensions by summing over the height and width $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each patch of the latent map.
@@ 3.1.2 Cosine Prototypes
which is based on the angle between two normalized vectors of unit length. By normalizing $D$ dimensional vectors to unit length, they are projected onto the surface of a $D$ dimensional hypershere. The cosine similarity is defined on the interval [1,1][-1,1] and measures: $1$ for two vectors pointing in the same direction, 0 for orthogonal vectors, and 1-1 for vectors pointing in opposite directions. Without the normalization to unit length, we can define a scaled dot-product prototype inspired by the Attention mechanism (Vaswani et al., 2017). However, this similarity measure is not bounded to any interval, as the vector length is proportionally preserved. For more details see Appendix E. Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing two vectors directly.
@@ 3.2.1 Gaussian Prototypes
The term inside the exponent is based on the $L_{2}$ distance between the latent vector $\bm{z}$} and the mean $\bm{{\mu}}$. The PDF formulation has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space. B) Prototypes can adapt their shape using a full covariance matrix, allowing different variances along various feature dimensions, offering more flexibility in shaping the latent space. However, learning a full covariance matrix increases computational requirements.
@@
Figure 3: HyperPG prototypes learn a distribution of cosine similarities. They use a learnable anchor vector $\bm{{\alpha}}$, scalar mean $\mu$ and variance $\sigma^{2}$. They project a Gaussian distribution of cosine similarities on the surface of a hypersphere, resulting in ring shaped activation patterns around the anchor vector.
@@ 3.2.2 Gaussian Prototypes on the Hypersphere - HyperPG
Just as Gaussian prototypes provide a probabilistic formulation for Euclidean prototypes, we aim to develop a probabilistic formulation for hyperspherical prototypes based on the cosine similarity. We name this new formulation Prototypical Gaussians on the Hypersphere (HyperPG). HyperPG prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$. HyperPG prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$. Because the cosine similarity is bounded to [1,1][-1,1], HyperPG’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds. Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function. Then, HyperPG’s similarity measure based on the truncated Gaussian distribution is defined as
@@
Figure 3 illustrates the activations of HyperPG’s similarity function on the surface of a $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, fixed std $\sigma=0.1$ and various mean values $\mu\in[0,1]$. The anchor $\bm{{\alpha}}$ defines a prototypical direction vector in latent space $\mathcal{Z}$, similar to other hyperspherical prototypes, and is visualized as a red arrow. The learned Gaussian distribution of cosine similarities is projected onto the hypersphere’s surface with std $\sigma$ governing the spread of the distribution, and mean $\mu$ the expected distance to the anchor $\bm{{\alpha}}$. For $\mu=1$, the distribution centers around the anchor, as the cosine similarity is $1$ if two vectors point in the same direction. This corresponds to the von Mises-Fisher (vMF) distribution (Appendix D). For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is 1-1 for vectors pointing in opposite directions.
For values of $1>\mu>-1$, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$. This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction. Imagining the interpolation between $\mu=1$ and $\mu=-1$, the probability mass moves from one pole of the hypersphere to the other, stretching like a rubber band over the surface. For $\mu=0$, the expected cosine similarity indicates that vectors with the highest activation are orthogonal to the anchor $\bm{{\alpha}}$. Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation.
@@
One drawback of the HyperPG formulation, is that it cannot be “pushed” to the nearest training sample as with the other formulations. However, the prototype’s interpretability can be maintained by following the approach of prototypical concept balls (Ma et al., 2023). Instead of only using the pushed training sample to interpret the prototype, all training sample with similar activation are equally good explanations. By presenting multiple sample patches the user is able to infer a learned concept.
HyperPG can be easily adapted to other probability distributions with additional desirable properties. Possible candidate distributions are elaborated on in Appendix D. Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPG idea transferable to other manifolds and applications outside of prototype learning.
4 Training Prototypical Part Networks
Prototypical Part networks like ProtoPNet have a reputation for being difficult to tune and hyperparameter-sensitive. To better explore the design choices, we categorize them into three groups: (1) Network architecture governs information flow from input to the output, including pretrained backbone network and neck for feature extraction. (2) Prototype hyperparameters shape the prototypical latent space through the choice of prototype formulation, but also number of prototypes, prototype dimensions and additional shaping losses. (3) Network optimization settings such as warmup-epochs, learning rate schedules or number of optimizers. ProtoPNet and its successors often rely on complex optimization schemes with warm-up phases and stage-specific optimizers and schedules. Detailed hyperparameter values are provided in subsection B.2.
4.1 Network Architecture
For our experiments, we use the ProtoPNet’s architecture (Chen et al., 2019) as the foundational deep prototype learning model for image classification. Figure 4 illustrates this network architecture with the HyperPG prototype module. The encoding stage employs a pretrained backbone (e.g., ResNet (He et al., 2016)) . A subsequent projection neck consisting of two $1\times 1$ convolution layers with ReLU between them and Sigmoid activation at the end, reduce dimensionality to create the latent prototype space.
In the prototype stage, similarity scores are computed. Original ProtoPNet uses Euclidean prototypes measuring $L_{2}$ distance between encoded features and prototype parameters. The HyperPG module, shown in Figure 4, first calculates cosine similarity between learnable anchors $\bm{{\alpha}}$ and neck-produced latent vectors. Then, a Density Estimation layer with parameters $\mu$ and $\sigma$ computes Gaussian PDF over the previous layer’s activations. Both components could be independently modified in future work, implementing alternative similarity measures (e.g., hyperbolic) or multi-modal probability distributions.
The classification stage produces output logits from similarity scores through a single linear layer, preserving interpretability.
4.1.1 Comparison to Recent Works
Our experiments use the basic ProtoPNet architecture, focusing on the effect of different prototype formulations. This section briefly outlines modifications required to implement the models in Table 1. LucidPPN (Pach et al., 2024) replaces the encoding stage with a hybrid encoder for color-only and texture-only feature extraction. Deformable ProtoPNet (Donnelly et al., 2021) modifies the prototype module’s input to allow dynamic spatial arrangements (“deformations”). ProtoPool (Rymarczyk et al., 2022) alters the prototype module’s output by learning to dynamically pool prototype activations. ProtoTree (Nauta et al., 2021) and ProtoKNN (Ukai et al., 2023) use a different inherently interpretable models for the classification stage. ProtoPFormer (Xue et al., 2024) utilizes a Vision Transformer backbone with separate network branches for global information (class token) and local information (image tokens).
4.2 Prototype Losses